Find the integral of (y2 -1 ) dx-2dy = 0


1
Expert's answer
2021-04-11T14:03:22-0400

"\mathrm{We\;have\;:\;\;}(y^2-1)\;dx-2\;dy=0\\ \;\\
\implies \;\dfrac{2\;dy}{y^2-1}=\;dx\\ \; \\
\textup{Integrating both sides , we get}\\ \; \\
\int \dfrac{2\;dy}{y^2-1}=\int dx\\ \; \\
\implies \;\int \dfrac{(y+1)-(y-1)}{(y+1)(y-1)}\;dy=\int dx"


"\implies \;\int \dfrac{1}{y-1}\;dy-\int \dfrac{1}{y+1}\;dy=\int dx\\ \;\\
\implies \;\ln |{y-1}|\;-\ln |{y+1}|=x+C\\ \;\\
\implies \;\ln \left|\dfrac{y-1}{y+1}\right|=x+C\\ \;\\
\implies \;\left|\dfrac{y-1}{y+1}\right|=Ae^x\\ \;\\"